OpenStudy (anonymous):

Find the distance from P to l. Line l contains points (6,5) and (2,3). Point P has coordinates (2,6).

OpenStudy (gamer56):

ok so the way to do this is like this The brute-force way to do this is: Find the slope of Line_I :   Mɪ = (5 – 3) ⁄ (6 – 2) = +½ The equation of Line_I is :      y = mx + b      y = (½)x + b ... substitute either point on Line_I      5 = (½)(6) + b      b = 2      y = (½)x + 2 ... Line_I equation The slope of the perpendicular (shortest) line from point_P to Line_I = Mpɪ  =  -1 ⁄ (Mɪ) = - 2  ——>  for Line_PI The equation of Line_PI is :  y = mx + b  y = - 2x + b ... substitute point_P  6 = - 2(2) + b  b = 10  y = - 2x + 10 ... Line_PI equation The intersection between Line_I and Line_PI is located at :         y = y     (½)x + 2 = - 2x + 10         x = 3.2 and the y_value at that x_location is:   y = (½)x + 2 ... Line_I equation   y = (½)(3.2) + 2   y = 3.6        y = - 2x + 10 ... Line_PI equation        y = - 2(3.2) + 10        y = 3.6  ...  checks out  ... so the intersection is at point_X = (3.2, 3.6)  ... and the distance from point_P to point_X is :    d = √  [ (∆x)² + (∆y)²  ]    d = √  [  (3.2 – 2)² + (6 – 3.6)²   ]    d = √  7.2  =  √  (36 ⁄ 5)  =  6(√ 5) ⁄ 5  =  2.68

OpenStudy (gamer56):

does this make since?

OpenStudy (anonymous):

Yesss. Thank you !!!!!